geerky42 (geerky42):

This is probably beyond my ability and knowledge, but I played around with math and I stumped on this: \[\Large \sum_{x=0}^{n-1}\left[\sum_{i=1}^k \binom{k}{i}x^{k-i}\right]=n^k\quad\quad \forall k\in\mathbb N\] I am convinced this is true, but how can one proves that? What do I need to know to prove it?

geerky42 (geerky42):

It is true for first 5; \[\sum_{x=0}^{n-1}1=n\]\[\sum_{x=0}^{n-1}2x+1=n^2\]\[\sum_{x=0}^{n-1}3x^2+3x+1=n^3\]\[\sum_{x=0}^{n-1}4x^3+6x^2+4x+1=n^4\]\[\sum_{x=0}^{n-1}5x^4+10x^3+10x^2+5x+1=n^5\] But of course this is not good enough to prove that this is true for all \(k\in\mathbb N\)...

OpenStudy (unklerhaukus):

induction?

OpenStudy (anonymous):

I am reminded of binomial theorem for the inner part.

ganeshie8 (ganeshie8):

I feel a combinatorics proof is possible as n^k reminds me of number of different strings of length "k" using "n" characters with repetition

OpenStudy (isaiah.feynman):

Yup. Binomial theorem here.

OpenStudy (anonymous):

\[ \sum_{i=1}^k \binom{k}{i}x^{k-i} = \left(\sum_{i=0}^k \binom{k}{i}x^{k-i}\right) - x^k = (1+x)^k-x^k \]

ganeshie8 (ganeshie8):

Yeah that telescopes nicely

geerky42 (geerky42):

lol just figured that out. Such beauty.

geerky42 (geerky42):

Pretty simply, thanks to Binomial Theorem. Thanks, everyone!